### v1

# Learning What to Write: Write-Gated KV for Efficient Long-Context Inference

Abstract
[…]
1 Introduction
The capability of Large Language Models (LLMs) to process extensive contexts has revolutionized applications ranging from long-document summarization (longbenchv2), repository-level code analysis (coderepoqa), to long-term agentic planning (planandact). However, this capability comes at a steep computational cost. While the model parameters are static, the Key-Value (KV) cache grows linearly with sequence length, creating a significant bottleneck for inference. In long-context scenarios, the sheer volume of the KV cache not only exhausts limited GPU memory but also saturates memory bandwidth, making the decoding process memory-bound. Consequently, the latency of generating each token increases proportionally with the context length, rendering real-time long-context interaction prohibitively expensive.
To mitigate this bottleneck, recent research has largely focused on two management strategies: KV Selection and KV Eviction. Selection-based methods, such as Quest (quest) and SparQ (sparq), selectively attend to relevant KVs at runtime without reducing the resident memory footprint. Eviction-based methods, such as H2O (h2o) and TOVA (tova), retrospectively prune the cache by identifying and removing unneeded tokens after they have been stored. While effective, these paradigms share a common, fundamental inefficiency: all generated tokens are indiscriminately committed to persistent memory. This results in “indiscriminate writing,” where the system blindly consumes memory resources to persist redundant or low-utility states, only to later skip or evict them during computation.
In this paper, we argue that an efficient long-context system requires a third, missing primitive: KV Admission. Instead of filtering information at read-time (Selection) or managing memory post-write (Eviction), KV Admission operates pre-write, acting as a gatekeeper that predicts the future utility of a token before it enters the KV cache.
To realize this, we introduce Write-Gated KV (WG-KV), a novel, learnable KV Admission mechanism. Unlike heuristic approaches that rely on fixed patterns, WG-KV employs a lightweight, differentiable predictor that evaluates token utility before cache entry. During prefilling, this predictor constructs a Vertical-Slash attention mask, ensuring that high-utility tokens are visible to the entire sequence while others are accessible only locally. Simultaneously, it routes generated KV pairs to distinct memory regions: tokens predicted as high-utility are admitted to a persistent Global Cache, while others are confined to a transient, sliding Local Cache. This selective persistence strategy seamlessly extends to the decoding phase, where we maintain this dual-cache architecture to filter out low-utility states at the point of KV entry.
[…]
Our key contributions are summarized as follows:
  • We formalize the distinction between Admission, Selection, and Eviction, identifying Admission as a critical missing primitive for efficient inference.
  • We propose WG-KV, a learnable admission mechanism that achieves high sparsity via a lightweight predictor.
  • We provide a hardware-aware realization of WG-KV that integrates with sparse FlashAttention (minference) and PagedAttention (pagedattention) kernels, ensuring that theoretical sparsity translates to wall-clock speedups.
  • We show that WG-KV provides substantial efficiency gains while being complementary with existing KV Selection methods.
2 Background and Motivation
2.1 The Attention Cost in Long-Context Inference
The fundamental building block of the Transformer architecture (attention) is the Scaled Dot-Product Attention. Given a sequence of input tokens, the attention output for a query vector $q_{t}\in\mathbb{R}^{d}$ at step $t$ is computed by attending to the historical Key-Value (KV) pairs $\{(k_{i},v_{i})\}_{i=1}^{t}$. Specifically, let $K_{\leq t}\in\mathbb{R}^{t\times d}$ and $V_{\leq t}\in\mathbb{R}^{t\times d}$ denote the matrices formed by stacking the key and value vectors up to the current step. The attention operation is defined as:
$\text{Attn}(q_{t},K_{\leq t},V_{\leq t})=\text{softmax}\left(\frac{q_{t}K_{\leq t}^{\top}}{\sqrt{d}}\right)V_{\leq t}$
In standard autoregressive generation, recomputing $K_{\leq t}$ and $V_{\leq t}$ from all previous tokens at every step would be computationally prohibitive. To address this, inference engines employ a KV Cache, storing the key and value vectors for previous tokens. At each step $t$, the system only computes the new pair $(k_{t},v_{t})$, appends it to the cache, and performs the attention operation over the accumulated state. While this caching mechanism is essential for autoregressive generation, the attention operation presents distinct bottlenecks during the two phases of inference: prefilling and decoding.
Figure 1: Attention Bottleneck in Long-Context Inference. Measured on Llama 3.1 8B (Batch Size=1) on an H200 GPU. As sequence length increases, the computational cost of attention dominates (a) prefill latency, while the overhead of the KV cache dominates both (b) decode latency and (c) memory usage.
During prefilling, the model processes the entire input prompt of sequence length $N$ in parallel. While highly parallelizable, the quadratic complexity of attention $O(N^{2})$ quickly dominates execution time. Consequently, attention becomes the primary bottleneck for long contexts. As illustrated in Figure 1a, for a sequence length of 100K tokens, attention computation accounts for 82% of the total latency, rising to 94% at 400K tokens.
During decoding, the model generates tokens autoregressively. Although the attention complexity per step is $O(N)$, the process is fundamentally memory-bound. Each generation step requires loading the KV cache from VRAM to GPU registers. As the sequence grows, the KV cache footprint expands linearly, eventually consuming the vast majority of GPU memory (Figure 1c). This linear growth in memory consumption not only increases latency (Figure 1b) but also hard-limits the maximum batch size and context length due to memory exhaustion.
2.2 A Causal View on KV Management
Table 1: Taxonomy of KV Cache Management Primitives.
Decision Scope Mechanism Inference Speed Memory Footprint Information Fidelity
KV Admission Pre-Write Predictive Filtering $\uparrow$ High $\downarrow$ Low $\approx$ High
(Ours) (Future Utility) Store only valuable KVs Small KV cache Compressed state Preserves semantics
KV Selection Read-Time Query-Guided Masking $\uparrow$ High == High $\approx$ High
(e.g., Quest, SparQ) (Current Query) Skip unrelated KVs Sparse attention Full state retained Full context access
KV Eviction Post-Write Retrospective Pruning $\uparrow$ High $\downarrow$ Low $\downarrow$ Risk of Degradation
(e.g., H2O, TOVA) (Past Statistics) Drop unneeded KVs Small KV cache Bounded state Information loss
Formally, we can model the KV cache $\mathcal{C}_{t}$ at time step $t$ as a set of key-value pairs stored in memory. In standard autoregressive generation, the cache management policy is trivial: it is an append-only operation where every new token’s state $(k_{t},v_{t})$ is permanently persisted.
$\mathcal{C}_{t}^{\text{std}}=\mathcal{C}_{t-1}^{\text{std}}\cup\{(k_{t},v_{t})\}$
This “indiscriminate writing” policy leads to linear memory growth $|\mathcal{C}_{t}^{\text{std}}|=t$ and becomes the root cause of attention bottleneck. To address this, prior works have introduced constraints on the attention operation or $\mathcal{C}_{t}$ itself. We unify these approaches under a causal framework based on the decision scope and timing of the operation, as summarized in Table 1.
KV Selection (Read-Time). Selection methods, such as Quest (quest) or SparQ (sparq), optimize the read path. They do not necessarily reduce the resident memory footprint of $\mathcal{C}_{t}$, but rather approximate the attention operation by selecting a subset of keys $\mathcal{S}_{t}\subset\mathcal{C}_{t}$ relevant to the current query $q_{t}$.
$\text{Attn}(q_{t},\mathcal{C}_{t})\approx\text{Attn}(q_{t},\text{Select}(q_{t},\mathcal{C}_{t}))$
While Selection reduces the computational cost from $O(t^{2})$ to $O(t\cdot k)$, where $k$ is the budget for selected tokens, the underlying cache $\mathcal{C}_{t}$ remains fully populated to support future queries, leaving the linear memory growth unresolved.
KV Eviction (Post-Write). Eviction methods, such as H2O (h2o) or TOVA (tova), impose a hard constraint $B$ on the cache size. These methods operate retrospectively: they first admit the new token into the cache, and then prune the set based on historical attention statistics (e.g., accumulated attention scores as in h2o).
$\mathcal{C}_{t}^{\text{new}}=\text{Evict}(\mathcal{C}_{t}^{\text{old}})\quad\text{s.t.}\quad|\mathcal{C}_{t}^{\text{new}}|\leq B$
Eviction effectively bounds memory usage. However, it suffers from a “write-then-throw” inefficiency: the system temporarily admits noise into the resident state, forcing subsequent decoding steps to waste memory bandwidth and computation attending to these tokens until they are eventually identified as redundant. Furthermore, eviction can be risky; information discarded is permanently lost, potentially degrading performance if the evicted tokens are needed for future queries (infinigen).
KV Admission (Pre-Write). We identify Admission as the missing third primitive. Unlike Eviction, which reacts to cache overflow, Admission acts as a gatekeeper. It evaluates the potential future utility of a new state $(k_{t},v_{t})$ before it is committed to $\mathcal{C}_{t}$.
$\mathcal{C}_{t}=\text{Admit}(\mathcal{C}_{t-1},k_{t},v_{t})$
Ideally, an Admission policy filters out noise tokens immediately, ensuring that $\mathcal{C}_{t}$ contains only high-utility information without ever paying the cost of writing redundant states. This offers a unique advantage: by preventing cache pollution at the source, Admission reduces memory growth and alleviates the memory access bottleneck of subsequent decoding.
Figure 2: (a) Synergy of KV Admission and Selection. (b) Transient Utility of Tokens. Visualization of attention scores received by a representative “low-utility” token over time. While the token is ignored by distant future queries, it remains highly relevant within a short local window.
A Unified View. These primitives are not mutually exclusive; rather, they operate at different stages of the token lifecycle. An ideal efficient inference system can employ Admission to filter the input stream, Selection to focus the attention computation, and Eviction to prune obsolete history from the admitted set. As illustrated in Figure 2(a), combining Admission with Selection enables the system to push the Pareto frontier of efficiency further than either method could achieve in isolation.
Figure 3: Heterogeneity of Token Utility. Attention patterns show that utility is both sparse and head-specific.
2.3 Sparsity and Locality in Attention Patterns
To design an effective KV Admission policy, we must first understand the characteristics of attention patterns in LLMs. We employ Llama 3.1 8B (llama3) to perform a long-context code summarization task using a sample from The Stack dataset (thestack) to observe which tokens are attended to and when. We identify three key properties:
Skewed Utility Distribution. Attention weights are not uniformly distributed; they exhibit extreme sparsity where a small subset of tokens attracts the vast majority of attention. As shown in Figure 3, specific tokens (e.g., Token 383 in Layer 10, Head 19) act as high-utility tokens, consistently receiving high attention scores from almost all future queries. In contrast, other tokens (e.g., Token 2196 in the same head) are virtually ignored after generation. This suggests that a large portion of the KV cache stores “noise” that contributes negligible value to future representations, validating the feasibility of aggressively filtering KVs via Admission.
Head-Specific Relevance. The utility of a token is not intrinsic to the token itself but is highly dependent on the attention head. A token that is critical for one head may be irrelevant for another. For instance, comparing the two heads in Figure 3, we observe that while Token 383 is a dominant anchor for Layer 10, Head 19, it is almost completely ignored by Layer 25, Head 3. Conversely, Token 2196, which was ignored by the former, becomes highly salient for the latter. This implies that a single unified admission policy (e.g., dropping a token from all layers simultaneously) is suboptimal. An effective Admission mechanism must be fine-grained, making independent write decisions for each head.
Figure 4: Memory Fragmentation with Ragged KV States.
The Necessity of Local Context. While long-term attention is sparse, short-term attention is dense. Figure 2(b) illustrates the attention score received by a “low-utility” token (Token 2472) over time. Although this token eventually receives near-zero attention from distant queries, it receives significant attention from its immediate successors. This “attention spike” in the recent window creates a dilemma for any admission policy: a token cannot be identified as “low utility” immediately upon generation, as it is temporarily high-utility. This necessitates a hybrid retention mechanism distinct from immediate filtering. Ideally, such a mechanism should guarantee a grace period for recent tokens to satisfy local attention density, before transitioning to a utility-based admission policy for sustaining long-range context.
2.4 Systems Challenges with Ragged KV States
Implementing a fine-grained admission policy introduces non-trivial systems challenges. In standard dense attention, the KV cache for a sequence is represented as a contiguous tensor. However, independent admission decisions by each head, as motivated in Section 2.3, result in a cache that is “ragged” across two dimensions:
  1. 1.
    Across Heads: Different heads prioritize different tokens (see Figure 3), leading to uneven cache lengths within the same layer.
  2. 2.
    Across Layers: The level of semantic abstraction varies by depth, causing sparsity rates to diverge between shallow and deep layers.
A naive implementation that simply concentrates admitted tokens leads to severe memory fragmentation, as shown in Figure 4. Pre-allocating maximum-length buffers negates the memory savings, while dynamic reallocation incurs prohibitive runtime overhead. To translate theoretical sparsity into practical wall-clock speedup and memory reduction, the system must leverage a decoupled logical-physical memory mapping, similar to PagedAttention (pagedattention), to manage these irregular, head-specific cache lifecycles efficiently.
3 Method
In this section, we present Write-Gated KV, a learnable KV Admission mechanism designed to eliminate the “indiscriminate writing” inefficiency in long-context inference. We first introduce the architectural overview of WG-KV (Section 3.1). Then, we derive the differentiable gating mechanism that enables end-to-end learning of token utility (Section 3.2). Finally, we formulate the training objective as a constrained optimization problem and present our relaxation strategy for practical training (Section 3.3).
3.1 Architectural Overview
The core philosophy of WG-KV is to decide what enters memory, filtering out noise at the source to prevent cache pollution. As illustrated in Figure 5a, our architecture introduces a lightweight predictor, the Write-Gate MLP, into the standard attention block.
For a given input sequence, standard LLMs indiscriminately store the key-value pairs $(k_{t},v_{t})$ for every token $t$. In contrast, WG-KV computes a scalar admission score $g_{t}\in[0,1]$ for each head, estimating the future utility of the token. Based on this score, the system routes the KV pair to one of two logical memory regions:
  • Global Cache: A persistent memory region for tokens with high predicted utility ($g_{t}\approx 1$). These tokens are retained persistently to support long-term dependencies.
  • Local Cache: A transient, fixed-size memory buffer for tokens with low predicted utility ($g_{t}\approx 0$). These tokens are discarded once they fall out of the recent context window.
This dual-cache design is motivated by the observation that even “low-utility” tokens often require high attention scores from their immediate successors (Section 2.3). By guaranteeing a sliding local window, we preserve the dense attention patterns typically observed in the recent window, while the learned gate $g_{t}$ aggressively sparsifies the long-range context.
3.2 Differentiable Gating Mechanism
To learn the admission policy, we integrate the gate $g$ into the attention formulation. We employ a lightweight Multi-Layer Perceptron (MLP) to compute $g$. We first define the input feature $x_{l,h,t}$ for layer $l$ head $h$ at step $t$ as the concatenation of the pre-RoPE key $k_{l,h,t}$ and post-RoPE key $k_{l,h,t}^{\text{rope}}$:
$x_{l,h,t}=[k_{l,h,t};k_{l,h,t}^{\text{rope}}]$
The admission score $g_{l,h,t}$ is then computed as:
$g_{l,h,t}=\sigma\left(W^{2}_{l,h}\cdot\text{GELU}(W^{1}_{l,h}\cdot x_{l,h,t}+b^{1}_{l,h})+b^{2}_{l,h}\right)$
where $W^{1}_{l,h},b^{1}_{l,h}$ and $W^{2}_{l,h},b^{2}_{l,h}$ are learnable projection matrices and bias vectors, and $\sigma$ is the sigmoid function.
Write-Gated Attention.
Standard Scaled Dot-Product Attention computes the output for a query $q_{i}\in\mathbb{R}^{d}$ against keys and values $\{(k_{j},v_{j})\}_{j\leq i}$ as:
$\text{Attn}(q_{i},K_{\leq i},V_{\leq i})=\sum_{j\leq i}\frac{\exp(q_{i}\cdot k_{j}/\sqrt{d})}{\sum_{m\leq i}\exp(q_{i}\cdot k_{m}/\sqrt{d})}\,v_{j}$
To simulate the effect of admission during training, we modulate the attention weights by the gate $g_{j}$. Critically, to preserve the recent context described in Section 3.1, we enforce a transient utility window of size $W_{\text{local}}$. Tokens within $W_{\text{local}}$ are always fully visible, while distant tokens are weighted by their predicted utility $g_{j}$:
$\tilde{g}_{ij}=\begin{cases}1&\text{if }i-j<W_{\text{local}}\\ g_{j}&\text{otherwise}\end{cases}$
Our proposed Write-Gated Attention is then formulated as:
$\widetilde{\text{Attn}}(q_{i},K_{\leq i},V_{\leq i})=\sum_{j\leq i}\frac{\exp(q_{i}\cdot k_{j}/\sqrt{d})\cdot\tilde{g}_{ij}}{\sum_{m\leq i}\exp(q_{i}\cdot k_{m}/\sqrt{d})\cdot\tilde{g}_{im}}\,v_{j}$
This formulation ensures that if the predictor outputs $g_{j}=0$, the token $j$ effectively vanishes from the attention of distant queries, mimicking the behavior of not writing it to the Global Cache. Conversely, if $g_{j}=1$, the token remains visible to future queries, simulating its admission to Global Cache.
Figure 5: Overview of Write-Gated KV.
Figure 6: Dual-Cache Paged Memory Management. (a–c) To handle ragged cache growth across heads, we decouple the logical view into a fixed-size Local Cache and a growing Global Cache. These are mapped to a unified physical KV Pool via per-head Page Tables to prevent memory fragmentation. (d) The cache update logic during decoding. We employ a Lazy Promotion strategy where a token is inspected upon replacement from the local ring buffer; it is promoted to the global cache only if its $g\geq\tau$.
Log-Space Transformation for Efficient Training.
While the multiplicative gating above is theoretically sound, standard attention kernels (flashattn) are optimized for the computation of $\exp(q\cdot k)$. Introducing a per-token multiplication factor disrupts the standard kernel optimizations. To address this, we transform the gating operation into the log-space bias:
$\exp\left(q_{i}\cdot k_{j}/\sqrt{d}\right)\cdot\tilde{g}_{ij}=\exp\left(\frac{q_{i}\cdot k_{j}}{\sqrt{d}}+\log(\tilde{g}_{ij})\right)$
This transformation allows us to reformulate Write-Gated Attention as:
$\widetilde{\text{Attn}}(q_{i},K_{\leq i},V_{\leq i})=\text{softmax}\left(\frac{q_{i}K_{\leq i}^{\top}}{\sqrt{d}}+B_{\text{gate}}\right)V_{\leq i}$
where $B_{\text{gate}}$ is an attention bias matrix derived from $\log(\tilde{g}_{ij})$. In practice, we compute $\log(\tilde{g}_{ij}+\epsilon)$ with small $\epsilon$ to ensure numerical stability. This log-space transformation allows us to leverage high-performance compiler-based attention kernels, such as FlexAttention (flexattn), enabling efficient training on long sequences without custom CUDA kernels.
3.3 Training Objective
We formulate the learning of the admission policy as a constrained optimization problem over the gating parameters $\theta=\{W^{1}_{l,h},b^{1}_{l,h},W^{2}_{l,h},b^{2}_{l,h}\}_{\forall l,h}$. Our goal is to minimize the degradation of the model’s representation capabilities while bounding the size of the KV cache.
Let $\mathcal{L}_{\text{distill}}(\theta)$ be the distillation loss measuring the divergence between the original model and the Write-Gated model, and let $\mathcal{M}(\theta)$ be the size of the KV cache. We aim to solve
$\min_{\theta}\mathcal{L}_{\text{distill}}(\theta)\quad\text{subject to}\quad\mathcal{M}(\theta)\leq B$
where $B$ is the target memory budget.
Lagrangian Relaxation.
Directly optimizing with this constrained problem is challenging. Following the method of Lagrangian multipliers, we relax the problem into an unconstrained objective:
$\min_{\theta}\mathcal{L}_{\text{distill}}(\theta)+\lambda\mathcal{M}(\theta)$
Here, $\lambda$ is a hyperparameter controlling the trade-off between accuracy and sparsity. We show in Appendix A that minimizing this weighted sum is equivalent to solving the constrained problem for a specific budget implicit to $\lambda$.
Differentiable Sparsity Proxy.
The cache size $\mathcal{M}(\theta)$ is discrete (counting admitted tokens), which is non-differentiable. To enable gradient-based learning, we introduce a soft sparsity loss $\mathcal{M}_{\text{soft}}(\theta)$ as a proxy. We define the per-token cost as g+g(1g)g+g(1-g), which penalizes the expected KV cache size while encouraging the gate to converge towards binary values (0 or 1). Averaged over $L$ layers, $H$ heads, and $T$ tokens, the sparsity loss is:
$\mathcal{M}_{\text{soft}}(\theta)=\frac{1}{L\cdot H\cdot T}\sum_{l,h,t}\Big(g_{l,h,t}+g_{l,h,t}(1-g_{l,h,t})\Big)$
This regularization term pushes the model to minimize the number of tokens written to the Global Cache. We combine this with a standard L2 distillation loss on the hidden states of the final layer to compose the total loss $\mathcal{L}_{\text{total}}(\theta)=\mathcal{L}_{\text{distill}}(\theta)+\lambda\mathcal{M}_{\text{soft}}(\theta)$. During inference, we apply a threshold $\tau$ (empirically set to 0.1) to discretize the continuous $g$ into binary decisions.
4 System Implementation
Implementing WG-KV presents a unique systems challenge: unlike standard Transformers where the KV cache grows uniformly across all heads, our admission policy results in a ragged cache structure with highly irregular lengths across heads (as discussed in Section 2.4). To translate theoretical sparsity into wall-clock speedup, we present a hardware-aware implementation that leverages paged memory management to handle irregular cache growth efficiently.
4.1 Dual-Cache Paged Memory Management
To address the memory fragmentation caused by head-specific admission policies, we decouple the logical view of the cache from its physical storage. We partition the logical KV cache of each attention head into two distinct components (Figure 6a) backed by a unified physical KV Pool (Figure 6b):
  1. 1.
    Local Cache: A fixed-size ring buffer of capacity $W_{\text{local}}$ storing the most recent tokens regardless of utility.
  2. 2.
    Global Cache: A dynamically growing buffer storing high-utility tokens ($g\geq\tau$). This region grows monotonically as new high-utility tokens are identified.
To bridge the logical and physical views, we maintain a per-head Page Table (Figure 6c). Physical memory is managed in fixed-size pages (set to 16 tokens in our implementation), and the page table maps the logical pages of the Local Cache and Global Cache to non-contiguous physical pages. This allows the Global Cache to grow dynamically without requiring contiguous reallocation.
4.2 Efficient Prefilling via Vertical-Slash Attention
During the prefill phase, the system processes the entire input sequence in parallel. The goal is to compute attention and populate the KV cache efficiently.
Vertical-Slash Sparse Attention. To accelerate attention computation, we leverage the sparsity pattern dictated by the gate value $g$. As shown in Figure 5b, the attention mask $M$ exhibits a ”Vertical-Slash” pattern: every token attends to globally high-utility tokens (“Vertical”) and its local neighborhood (“Slash”). Formally, for a query at position $i$ and a key at position $j$, the attention mask $M_{ij}$ is defined as:
$M_{ij}=\mathbb{I}(i-j<W_{local})\lor\mathbb{I}(g_{j}\geq\tau)$
where $\mathbb{I}$ is the indicator function. We utilize the Vertical-Slash sparse attention kernel from MInference (minference) to compute attention efficiently. This avoids the $O(N^{2})$ cost of full attention by skipping computations for masked regions.
Initial Cache Population. Concurrently with the attention computation, we write the KV states to memory. Tokens within the final $W_{\text{local}}$ window are written to the Local Cache. Tokens preceding this window are written to the Global Cache only if their gate score satisfies $g\geq\tau$; otherwise, they are discarded immediately. This ensures the decode phase begins with a compact state while aligning with the training objective described in Section 3.3.
4.3 Efficient Decoding via Lazy Promotion
In the decode phase, the system generates tokens autoregressively. Unlike standard inference which continuously appends new KVs to a monotonically growing cache, WG-KV manages the dual-cache structure efficiently while handling the irregular cache lengths. We implement a Lazy Promotion strategy that strictly enforces the local window size $W_{\text{local}}$ via a ring buffer while selectively persisting high-utility states.
Cache Update Logic. The update mechanism operates at the granularity of individual heads. As illustrated in Figure 6d, when a new token $(k_{t},v_{t})$ is generated with a predicted gate score $g_{t}$, the following operations are performed:
  1. 1.
    Inspection: The system inspects the existing KV pair at the current position of the Local Pointer. This pair represents the oldest token within the current local window and is the candidate to be overwritten.
  2. 2.
    Promotion: We check the gate score associated with this “victim” token. If its score satisfies the retention threshold ($g\geq\tau$), the token is identified as high-utility. It is then promoted to the Global Cache by copying its state to the next slot in the Global Cache and advancing the Global Pointer. This ensures that only valuable information is persisted beyond the local window. Tokens with $g<\tau$ are dropped silently.
  3. 3.
    Replacement: Finally, the new state $(k_{t},v_{t})$ and its gate score $g_{t}$ are written into the Local Cache at the current Local Pointer position, overwriting the old state. The Local Pointer is then incremented (modulo $W_{\text{local}}$).
Kernel Compatibility via Batch Transposition. Because the admission decisions are head-specific, the length of the Global Cache can vary significantly across heads (as shown in Figure 6a). To batch these irregular computations efficiently, we adapt the PagedAttention kernel (pagedattention). Standard PagedAttention supports variable sequence lengths across the batch dimension. We exploit this by reshaping our attention input: we fold the head dimension ($H$) into the batch dimension ($B$), treating the workload as $B\times H$ independent sequences of varying lengths. This allows us to utilize optimized CUDA kernels to compute attention over the ragged KV cache efficiently.
5 Experiments
In this section, we evaluate WG-KV across three key dimensions: (1) the trade-off between KV cache size and task accuracy on long-context benchmarks, (2) the practical system efficiency in terms of latency and memory reduction, and (3) the composability with existing read-time KV selection methods.
Figure 7: Long-Context Performance on HELMET. We compare WG-KV against static, input-independent admission baselines (Local Attention and DuoAttention) across 14 tasks with a 32k context length. The x-axis represents the normalized KV cache size. WG-KV maintains high accuracy even in the low-memory regime ($40\%$), whereas baselines degrade significantly as the cache size decreases.
5.1 Experimental Setup
[…]
Baselines. To rigorously evaluate the effectiveness of our learnable admission policy, we compare against two representative baselines which we re-contextualize as static, input-independent admission policies:
  • Local Attention. A uniform admission policy that only admits the most recent tokens while retaining a small set of initial sink tokens (streamingllm).
  • DuoAttention (duoattn). A head-wise static policy that designates specific heads as “retrieval heads” (full cache) and others as “streaming heads” (local cache only). We use the official configurations to determine the head allocation.
For all methods, we fix the local window size $W_{\text{local}}$ to 256 and sink size (if applicable) to 128 to ensure a fair comparison of short-term context retention.
Evaluation Benchmarks. We evaluate on HELMET (helmet), a comprehensive long-context benchmark. We report results on a subset of 14 tasks with a 32k context length, categorized into: Retrieval Augmented Generation (Natural Questions, TriviaQA, PopQA, HotpotQA), Passage Reranking (MS MACRO), Long-Document QA (NarrativeQA, Infinibench QA/MC), Summarization (Infinibench Sum, Multi-LexSum), and Many-Shot In-Context Learning (TREC Fine, NLU, Banking77, CLINC150).
5.2 Efficiency-Accuracy Tradeoff
We first investigate the efficacy of WG-KV in preserving model performance under constrained memory budgets. For WG-KV, we fix the binarization threshold $\tau=0.1$ while sweeping $\lambda$ to obtain different cache sizes. For baselines, we vary the sliding window size (Local Attention) or the ratio of retrieval heads (DuoAttention). Figure 7 presents the benchmark results.
WG-KV consistently outperforms both Local Attention and DuoAttention, particularly in the low-memory regime ($40\%$ cache size). Notably, in information-intensive tasks such as Passage Reranking (e.g., MS MACRO) and Summarization (e.g., InfiniteBench Sum), WG-KV maintains near-lossless performance even when reducing the KV cache to 10% of its original size. In contrast, Local Attention suffers from catastrophic degradation as soon as the context exceeds the local window, and DuoAttention, while more robust, fails to capture the dynamic, input-dependent utility of tokens, leading to sub-optimal admission. These results confirm our hypothesis that token utility is inherently dynamic and input-dependent; a learned gate that decides admission at runtime is significantly more effective than static heuristics.
Figure 8: System Efficiency and Hyperparameter Analysis. (a-c) End-to-end latency and memory usage on Llama 3.1 8B (H200 GPU) with 75% sparsity. WG-KV achieves up to $3.45\times$ prefill speedup and avoids OOM at 500k context. (d) Pareto frontier of the loss-efficiency trade-off on the validation set. We plot distillation loss against normalized cache size by sweeping $\lambda$ and $\tau$.
5.3 System Efficiency
To validate the practical efficiency of WG-KV, we measure end-to-end latency and memory usage on an NVIDIA H200 GPU. We simulate a 75% sparsity setting (i.e., retaining only 25% of the KV cache) and compare it against Full Attention. The measurements are performed with a batch size of 1 (single request). The results are shown in Figure 8a-c.
[…]
Memory Reduction. WG-KV reduces peak memory usage by 46–57% (Figure 8c). Crucially, at 500K context length, the Full Attention baseline triggers an Out-Of-Memory (OOM) error, whereas WG-KV successfully completes inference with reduced KV cache. This demonstrates that KV Admission is not just an optimization for speed, but also an enabler for processing ultra-long contexts.
Overhead Analysis. The introduction of the Write-Gate MLP incurs minimal overhead. In terms of model size, the additional MLP parameters constitute only $\sim$0.42% of the total parameter count in Llama 3.1 8B. In terms of computation, as indicated by the non-attention latency (blue bars) in Figure 8a-b, the latency overhead introduced by the predictor is negligible. Even at a batch size of 1, the massive reduction in attention computation and memory access latency far outweighs the marginal cost of the predictor.
Figure 9: Composability with KV Selection. We compare “Quest Only” (selection from full cache) against “WG-KV + Quest” (selection from a WG-KV compressed cache, $\lambda=0.08$) on HELMET tasks. The x-axis denotes the Quest selection budget (log scale). The nearly identical curves demonstrate that WG-KV’s write-time admission is complementary to Quest’s read-time selection.
5.4 Composability with KV Selection
A key conceptual contribution of our work is identifying KV Admission (write-time) and KV Selection (read-time) as distinct, complementary primitives. We validate this by combining WG-KV with Quest (quest), a representative KV Selection method. Figure 9 plots the benchmark results on HELMET tasks against the number of KV pairs attended to per query token. The “Quest Only” baseline performs selection on the full KV cache, while the “WG-KV + Quest” line applies Quest selection on top of a cache already compressed by WG-KV ($\lambda=0.08$). The curves are nearly identical, and in some cases, the combined approach marginally outperforms Quest alone. This result implies that WG-KV effectively filters out “noise” tokens that Quest would likely have disregarded anyway. By combining learning-based admission with query-guided selection, we achieve a compound reduction in computational cost without compromising information fidelity, as discussed in Section 2.2.
5.5 Hyperparameter Sensitivity
We analyze the impact of the training-time hyperparameter $\lambda$ and the inference-time binarization threshold $\tau$ on the trade-off between representational fidelity and memory usage. Figure 8d plots the distillation loss computed on a held-out validation set of FineWeb-Edu samples against the normalized KV cache size. By sweeping $\lambda$ and $\tau$, we trace the Pareto frontier of the model’s efficiency. Increasing $\lambda$ effectively drives the model towards smaller KV cache sizes, albeit with an increase in distillation loss. Empirically, we observe that fixing $\tau\approx 0.1$ consistently yields near-optimal operating points along this frontier across various $\lambda$ settings. Based on this observation, we adopt $\tau=0.1$ for all of our experiments.
6 Related Work
Static and Head-Level Attention Specialization.
Studies on attention dynamics reveal that token utility is structurally skewed across heads and positions. streamingllm identify “attention sinks,” demonstrating that retaining a few initial tokens allows for stable streaming inference without the full history. Building on this, fastgen propose profiling attention heads to statically categorize them for distinct retention policies. Mechanistic studies (retrievalhead) further reveal that specific “retrieval heads” are uniquely critical for long-range dependencies. Leveraging these characteristics, duoattention, razorattention, and prulong enhance efficiency by statically identifying retrieval-critical heads to retain full context, while restricting others to local context. Unlike these static heuristics, our approach introduces a learnable, input-dependent admission mechanism that captures dynamic utility patterns at runtime.
Read-Time KV Selection.
A large body of work accelerates long-context inference by selectively attending to a sparse, query-dependent subset of cached KV pairs at each step. These methods approximate attention while maintaining the complete underlying cache for future queries. Existing approaches exploit sparsity in various dimensions, including temporal (infllm; hip; specache), channel (sparq), inter-layer (omnikv), and low-rank (sea; loki) structures. Some retrieval-oriented designs treat the KV cache as an indexed memory (retrievalattn; squeezedattn; retroinfer; pqcache; clusterkv; arkvale; magicpig). Systems-oriented designs further integrate selection with KV-aware paging (quest; lserve), runtime management (infinigen; shadowkv), or learnable indexers (deepseekv32) to scale fine-grained selection to very long contexts. Our write-time admission mechanism is complementary to these read-time strategies, allowing for compound efficiency gains by filtering noise prior to selection.
Post-Write KV Eviction.
Another line of work focuses on maintaining a bounded memory by retrospectively pruning previously written states. Representative methods employ statistics derived from past attention or recency to evict low-utility tokens (h2o; tova; scissorhand; keyformer; sirllm; l2norm). Other approaches refine the eviction criteria through learnable interaction scores (dynamiccontextpruning; attentiongate), importance estimation (trimkv; locret), reconstruction objectives (kvzip), or prompt-guided heuristics (snapkv; nacl; pyramidinfer; finch). Unlike admission, which filters intrinsic noise at the source, eviction manages cache bounds by removing obsolete history; thus, the two serve distinct, complementary roles in the token lifecycle.
Representation-Level KV Compression.
A parallel line of research seeks to compress the KV representations themselves. Instead of discarding tokens, merging-based methods aggregate information by combining keys and values based on semantic similarity (kvmerger), weighted accumulation (cam), low-rank residual recovery (less), or learnable mechanisms (dmc). Additionally, dimensionality reduction techniques exploit inherent redundancy in hidden states by projecting KV pairs into lower-rank subspaces (eigenattn; palu; think; matryoshkakv), leveraging inter-layer redundancy (minicache; lckv; cla; xkv), or applying quantization (kvquant). These lines of work are largely orthogonal to our admission mechanism.
7 Conclusion
In this paper, we addressed the fundamental inefficiency of indiscriminate writing in long-context inference by formalizing KV Admission as a critical third primitive alongside Selection and Eviction. We introduced Write-Gated KV (WG-KV), a lightweight, learnable mechanism that predicts token utility at the point of generation. By aggressively filtering out noise before it enters memory, WG-KV achieves substantial reductions in memory usage and latency while preserving accuracy. Our results, particularly the composability with read-time selection methods, demonstrate that learning what to write, then reading only what is needed, constitutes a principled recipe for scalable inference.
Appendix A Theoretical Justification of the Optimization Objective
In Section 3.3, we formulated the training of Write-Gated KV as a constrained optimization problem, aiming to minimize the distillation loss $\mathcal{L}_{\text{distill}}(\theta)$ subject to a memory budget constraint $\mathcal{M}(\theta)\leq B$. To solve this, we employed a Lagrangian relaxation approach, minimizing the weighted sum $\mathcal{L}_{\text{total}}(\theta)=\mathcal{L}_{\text{distill}}(\theta)+\lambda\mathcal{M}(\theta)$.
In this appendix, we formally show that a solution to the relaxed unconstrained problem is effectively a solution to the original constrained problem for a specific budget implicit to $\lambda$.
A.1 Equivalence of the Relaxed Objective
Let $\theta$ denote the model parameters. We define the two objective functions as:
  • Unconstrained (Relaxed) Problem:
    $\theta^{*}(\lambda)=\operatorname*{argmin}_{\theta}\left[\mathcal{L}_{\text{distill}}(\theta)+\lambda\mathcal{M}(\theta)\right]$ (1)
  • Constrained Problem:
    $\min_{\theta}\mathcal{L}_{\text{distill}}(\theta)\quad\text{subject to}\quad\mathcal{M}(\theta)\leq B$ (2)
Proposition A.1.
For any $\lambda>0$, let $\theta^{*}(\lambda)$ be a global minimizer of the unconstrained problem defined in Eq. (1). Then, $\theta^{*}(\lambda)$ is also a global minimizer of the constrained problem defined in Eq. (2) where the budget is set to $B=\mathcal{M}(\theta^{*}(\lambda))$.
Proof.
By the definition of the global minimizer $\theta^{*}(\lambda)$, for any parameter setting $\theta$, the following inequality holds:
$\mathcal{L}_{\text{distill}}(\theta^{*}(\lambda))+\lambda\mathcal{M}(\theta^{*}(\lambda))\leq\mathcal{L}_{\text{distill}}(\theta)+\lambda\mathcal{M}(\theta)$ (3)
Now, consider any feasible solution $\theta$ for the constrained problem where the budget is $B=\mathcal{M}(\theta^{*}(\lambda))$. That is, consider any $\theta$ such that:
$\mathcal{M}(\theta)\leq\mathcal{M}(\theta^{*}(\lambda))$ (4)
We can rearrange the inequality from the unconstrained objective as:
$\mathcal{L}_{\text{distill}}(\theta^{*}(\lambda))-\mathcal{L}_{\text{distill}}(\theta)\leq\lambda\left(\mathcal{M}(\theta)-\mathcal{M}(\theta^{*}(\lambda))\right)$ (5)
Since $\lambda>0$ and we have restricted our consideration to $\theta$ where $\mathcal{M}(\theta)-\mathcal{M}(\theta^{*}(\lambda))\leq 0$, the right-hand side of the inequality is non-positive:
$\lambda\left(\mathcal{M}(\theta)-\mathcal{M}(\theta^{*}(\lambda))\right)\leq 0$ (6)
Therefore, it follows that:
$\mathcal{L}_{\text{distill}}(\theta^{*}(\lambda))-\mathcal{L}_{\text{distill}}(\theta)\leq 0\implies\mathcal{L}_{\text{distill}}(\theta^{*}(\lambda))\leq\mathcal{L}_{\text{distill}}(\theta)$ (7)
This demonstrates that among all parameters satisfying the memory constraint $\mathcal{M}(\theta)\leq\mathcal{M}(\theta^{*}(\lambda))$, the solution $\theta^{*}(\lambda)$ achieves the minimal distillation loss. Thus, $\theta^{*}(\lambda)$ is the optimal solution to the constrained problem for the specific budget $B=\mathcal{M}(\theta^{*}(\lambda))$. ∎
A.2 Discussion on Non-Convexity and Duality Gap
The proposition above guarantees that every solution found by sweeping $\lambda$ lies on the Pareto frontier of the accuracy-efficiency trade-off. Specifically, sweeping $\lambda$ traces the lower convex hull of the feasible region in the (Memory, Loss) plane.
However, deep neural network optimization landscapes are inherently non-convex. Consequently, Strong Duality may not hold, and there may exist a “Duality Gap.” This implies that for certain arbitrary budgets $B^{\prime}$, the true optimal solution might lie in a non-convex region of the Pareto frontier that cannot be reached by linear scalarization (i.e., minimizing $\mathcal{L}+\lambda\mathcal{M}$).
In the context of Write-Gated KV, this theoretical limitation is practically negligible. Our goal is not to target a specific, hard-coded memory budget (e.g., exactly 4096 MB), but rather to discover high-efficiency operating points where the marginal reduction in memory justifies the marginal cost in accuracy. By varying $\lambda$ (as performed in our experiments), we effectively uncover the set of Pareto-optimal solutions that form the convex hull of the trade-off curve, which represents the most resource-efficient configurations for deployment.